### v1

@@ Abstract
We present a machine learning framework to facilitate the solution of nonlinear multiscale differential equations and, especially, inverse problems using Physics-Informed Neural Networks (PINNs). This framework is based on what is called multihead (MH) training, which involves training the network to learn a general space of all solutions for a given set of equations with certain variability, rather than learning a specific solution of the system. This setup is used with a second novel technique that we call Unimodular Regularization (UR) of the latent space of solutions. We show that the multihead approach, combined with the regularization, significantly improves the efficiency of PINNs by facilitating the transfer learning process thereby enabling the finding of solutions for nonlinear, coupled, and multiscale differential equations.
@@
AI, PINNs, Transfer Learning, Multihead, Differential Geometry
@@ 1 Introduction
Non-linear differential equations describe complex systems in nature, such as fluid dynamics, climate modeling, and general relativity, to mention a few. While linear equations are widely applicable, non-linear equations often arise in systems where interactions or feedback loops create more intricate and unpredictable behaviors. Yet, in most cases, analytic solutions are either impossible or highly complex, necessitating the use of numerical methods to provide solutions [1]. With the advent of deep neural networks, particularly Physics-Informed Neural Networks (PINNs), it has become increasingly feasible to solve complex differential equations more efficiently. Extensive literature now exists on how to use PINNs to solve a wide variety of nonlinear, multiscale differential equations including those in complex, real-world systems where traditional numerical methods may struggle. An added layer of complexity arises when the goal is to solve inverse problems, where the objective is to infer unknown parameters or functions within a system based on available data, such as initial or boundary conditions [2].
Here, we present two efficient algorithms to use in PINNs that are aimed at significantly improving the solution of inverse problems for non-linear, coupled and multi-scale differential equations (DEs). The first technique is called multi-head training, where we train the network to learn the general space of all solutions (or latent space) to different variations of the same DEs, from which we can perform transfer learning to different regimes [3, 4, 5, 6, 7, 8]. The second one is a novel techinique that we call unimodular regularization of the latent space of solutions. It is based on the computation of the metric tensor, a concept from differential geometry, induced on the latent-space-hypersurface. This second technique facilitates (or even unlocks) the transfer learning process to extreme regimes by ensuring a more controlled response of the latent space to changes in the DEs [9].
There are previous works where concepts of differential geometry have been used in the context of NNs. In [10] the authors explore and characterize the behavior of NNs during training through the theory of Riemannian geometry and curvature - they relate the output of the NN with the input transformations through the curvature equations. An analysis and comparison between the convergence properties of the metric of NNs (used for classification) and the metric under the Ricci-DeTurck flow is done in [11]. In [12, 13], the authors remove edges of low importance (prune) the computational graph using the definition of Ricci curvature. To the best of our knowledge, differential geometry techniques have not been used in order to increase the performance of PINNs.
@@
Using PINNs to solve DEs offers several advantages over traditional methods. For instance, a PINN can be trained on a specific range of boundary conditions (BCs) or initial conditions (ICs). The specific solution(s) to a DE can only be determined if the corresponding ICs or BCs are specified as part of the problem. The existence and uniqueness theorem [14] ensures that there is a unique solution for a set of DEs once specific ICs or BCs are chosen. In traditional numerical methods, numerical integration is performed for each set of initial and boundary value conditions (IC-BV). If multiple IC-BVs are given, the numerical integration must be run separately for each one. However, for PINNs, we can train the NN on a specific set of DEs in a given range of IC-BVs referred to as a bundle, which must be sampled [15]. The training process may take more computational resources. However, once trained, the PINN can be evaluated almost instantaneously [16, 17] for any choice of IC-BV within the bundle, even if it was not trained specifically on it, and will still yield a solution. This may be extremely useful in cases where a very large number of solutions is required, each corresponding to different initial and boundary value conditions (IC-BV) or parameters [18].
@@
Stiff DEs, e.g. DEs where there is a large separation of scales between different features of the solution, can be challenging to solve [19]. Stiffness can manifest in various forms, including eigenvalue-based stiffness (where the system’s Jacobian has widely varying eigenvalues), multi-scale stiffness (involving both fast and slow time scales such as the flame equation [^1] Strictly speaking, the flame equation exhibits multiple types of stiffness. in Sec. 3.1), damping-induced stiffness (common in physical systems with strong damping such as the Van der Pol equation Sec 3.2), nonlinear stiffness, interval-dependent stiffness, and transient stiffness [20]. In the cases we present, traditional numerical methods are capable of solving the equations. However, in the context of solving DE systems with PINNs, we present novel techniques that significantly improve, or even unlock, our ability to solve them [21]. Moreover, these techniques are particularly relevant when there is no established numerical setup to solve the problem. This is the case of inverse problems. PINNs are a well-suited tool for solving such problems [17], as we will explain through the example of Sec. 3.3 (for a thorough analysis of this example see [22]).
@@
The purpose of this article is to introduce these techniques and show their efficiency with several examples. The algorithms that we present here are of general use in solving differential equations and inverse problems. The article is structured as follows: in section 2 we introduce the methodology used; we describe the NN structure in 2.1 and 2.2, the general training procedure in 2.3, and unimodular regularization in 2.4. The application of these techniques to different ODE systems and results are presented in section 3; we solve the Flame Equation in 3.1, the Van der Pol Oscillator in 3.2, and Einstein Field Equations in 5-dimensional Anti de Sitter spacetime in 3.3. The conclusions are summarized in section 4.
2 Methodology
@@
Physics-Informed Neural Networks (PINNs), first introduced in the works of Dissanayake and Phan-Thien [23] and Lagaris et al. [24], have emerged as a powerful framework for solving partial and ordinary differential equations (PDEs/ODEs). Raissi et al. [21] popularized this approach by demonstrating its effectiveness on a variety of challenging physical problems, while other researchers have made significant advancements in applying neural networks to PDEs and ODEs, including notable examples such as [25, 26, 27]. In this paradigm, one typically trains a distinct neural network for each unknown function in the governing equations. The training process involves minimizing a loss function that encodes the squared residuals of these differential equations, thereby embedding the physics directly into the learning objective.
@@
While traditional numerical methods often outperform PINNs in terms of computational efficiency and accuracy, PINNs offer unique advantages. One key benefit is their ability to learn the parametric dependence of solutions with respect to varying equation forms or initial/boundary conditions (IC/BC). In contrast, traditional methods typically require rerunning the solver for each set of different IC/BC. With PINNs, however, it is possible to train the network to simultaneously accommodate multiple conditions, enabling rapid evaluation of solutions for various IC/BC values or even different, parametrically related equations. This allows us to attack otherwise untreatable problems, like recovering a free form function appearing on the equations from different initial/boundary data of the uknown functions. Another notable advantage of using PINNs is their ability to leverage transfer learning, as discussed in Sec. 2.3. Transfer learning allows the model to generalize beyond the range of its training data, enabling the prediction of solutions in unseen regions or conditions. This technique has proven crucial for obtaining solutions in the stiffer regimes of certain differential equations [8]. Throughout this paper, we will leverage this concept within the so-called Multi-Head (MH) framework.
@@
The latent space of solutions can then be “projected down” to a specific solution for a given element of the family of the DE system by non-linearly combining the $d$ components. This projection is performed by another NN, known as the head. We will describe the multi-head setup in detail in the following sections. For pedagogical clarity, we will first discuss a single-head model to explain its implementation within the PINNs framework. Subsequent subsections will show how this approach can be generalized to multiple heads and how the latent space can be learned to extrapolate solutions to regions where the body has not been trained— a process known as transfer learning. Finally, we will introduce a novel technique called Unimodular Regularization, which improves the performance of the MH approach when transferring solutions to stiff regimes of differential equations.
@@ 2.1 Single-head
Figure 1: Scheme of the single head approach - we can see how this simplified case is the standard case for a deep neural network with multiple hidden layers and one single output. The output solutions are calculated as $\psi^{\text{NN}}=\texttt{head}_{W}\left(H\right)$ with $\psi\in\mathbb{R}^{(m\times 1)}$, $H\in\mathbb{R}^{(m\times d)}$ and $W\in\mathbb{R}^{(d\times 1)}$, being $m$ the batch size (number of points sampled in the independent variable $x^{1}$ meshed with sampling of $x^{2,3,\dots}$), and $d$ the number of neurons in the last layer of the head. The loss $L$ is of the form in Eq. (3). The red arrow represents the updating of the NN-parameters $w_{j}$ (weighs and biases) based on minimization of the loss function through gradient descent (represented by $\nabla_{w_{j}}L$).
@@
The sum over $\mu$ represents the total over the different elements of the DE system, including also all sampled points along the independent variable $x^{0}$. Finally, the sum [^2] Note that the distinction between summing and computing the mean is irrelevant, as it merely scales the loss by a factor inversely proportional to the number of sampled points. over the equations is the total of the residuals for each equation in our system of DEs.
@@ 2.2 Multi-head
The goal is to find a specific solution $\psi_{\alpha}^{\text{NN}}(x^{\mu})$ for each problem by projecting $H$ into a reduced space, non-linearly combining its components through a head with parameters $W_{\alpha}$, where the index $\alpha$ labels different variations of the DE family. Each head $W_{\alpha}$ correspond to a specific DE variation and produces a loss $L_{\alpha}$ as described in Eq. 3. We then define the total loss at each iteration as follows:
$L_{\text{tot}}=\sum_{\alpha=1}^{\#\text{heads}}L_{\alpha}$ (4)
@@
Figure 2: Scheme of the multi-head approach. The index $\alpha$ in the inputs $x^{\mu}_{\alpha}$ denotes all the variations within the DE family - all of them are inputs to the body, but each specific element of the family is later solved by its corresponding head, giving solutions $\psi_{\alpha}(x^{\mu}_{\alpha})$ as output ($\alpha=1,\dots,h$). The total loss $L_{\text{tot}}$ is as defined in (4). The backpropagation of the derivatives $\nabla_{w_{j}}L_{\text{tot}}$ indicated by the red arrow updates all the weights and biases $w_{j}$ (both of the body and all heads).
@@ 2.3 Training and transferring
Once the latent space $H$ stabilizes, we freeze the parameters of the body [^3] Freezing the parameters of the body implies detaching the tensors of $H$., so they no longer change, even as the model continues training. We denote the frozen body as $H^{f}$.
@@
The process of transfer learning (TL) can be used to interpolate, i.e. to obtain solutions for new elements of the DE family given by choices of parameters/IC/BC/forcing functions whose values fall within the range of training of the body $H^{f}$. TL can also be used for extrapolation; it can be used to find a specific solution to a variation within the DE family, even if that variation lies outside the values used in training. This is often the case when transferring to stiffer regimes. Next, we present a new technique used during training and show that it improves results during the transfer learning phase.
Figure 3: Scheme of the transfer learning procedure from a pre-trained body through the MH approach of section 2.2. The blue, diagonal lines in the body indicate that it is frozen, i.e. its weights and biases do not change. Only the weights of the new head are learnt in order to find the solution to a new variation of the DE system, as indicated by the red arrow. As explained in the text, this method can be used both for interpolation and extrapolation of solutions from the pretrained latent space.

### v2

@@ Abstract
Non-linear differential equations are a fundamental tool to describe different phenomena in nature. However, we still lack a well-established method to tackle stiff differential equations. Here we present a machine learning framework to facilitate the solution of nonlinear multiscale differential equations and, especially, inverse problems using Physics-Informed Neural Networks (PINNs). This framework is based on what is called multi-head (MH) training, which involves training the network to learn a general space of all solutions for a given set of equations with certain variability, rather than learning a specific solution of the system. This setup is used with a second novel technique that we call Unimodular Regularization (UR) of the latent space of solutions. We show that the multi-head approach, combined with Unimodular Regularization, significantly improves the efficiency of PINNs by facilitating the transfer learning process thereby enabling the finding of solutions for nonlinear, coupled, and multiscale differential equations.
@@
AI, PINNs, Transfer Learning, Multi-head, Differential Geometry
@@ 1 Introduction
Non-linear differential equations (DEs) describe complex systems in nature, such as fluid dynamics, climate modeling, and general relativity, to mention a few. While linear equations are widely applicable, non-linear equations often arise in systems where interactions or feedback loops create more intricate and unpredictable behaviors. Yet, in most cases, analytic solutions are either impossible or highly complex, necessitating the use of numerical methods to provide solutions [1]. With the advent of deep neural networks, particularly Physics-Informed Neural Networks (PINNs), it has become increasingly feasible to solve complex DEs more efficiently. Extensive literature now exists on how to use PINNs to solve a wide variety of nonlinear, multiscale DEs including those in complex, real-world systems where traditional numerical methods may struggle. An added layer of complexity arises when the goal is to solve inverse problems, where the objective is to infer unknown parameters or functions within a system based on available data, such as initial or boundary conditions [2].
@@
Using PINNs to solve DEs offers several advantages over traditional methods. For instance, a PINN can be trained on a specific range of boundary conditions (BCs) or initial conditions (ICs). The specific solution(s) to a DE can only be determined if the corresponding ICs or BCs are specified as part of the problem. The existence and uniqueness theorem [3] ensures that there is a unique solution for a set of DEs once specific ICs or BCs are chosen. In traditional numerical methods, numerical integration is performed for each set of ICs-BCs. If multiple IC-BVs are given, the numerical integration must be run separately for each one. However, for PINNs, we can train the neural network (NN) on a specific set of DEs in a given range of IC-BVs referred to as a bundle, which must be sampled [4]. The training process may take more computational resources. However, once trained, the PINN can be evaluated almost instantaneously [5, 6] for any choice of IC-BV within the bundle, even if it was not trained specifically on it, and will still yield a solution. This may be extremely useful in cases where a very large number of solutions is required, each corresponding to different ICs-BCs or parameters [7].
Previous work has been done on using NNs for learning the operator mapping free functions $u(x)$, BCs or ICs given as inputs that appear explicitly in the DEs, to the corresponding solutions $s(x)$. The operators learned in DeepONets [8] and Fourier Neural Operators (FNO) [9] can be used to predict new solutions $\tilde{s}(x)$ for stiffer input functions $\tilde{u}(x)$ that model has not trained on. In this sense, the end goal is similar to the one in this work: learning a higher-dimensional “response” of the NN (in our case, the latent space) that allows to compute solutions to different DEs by varying $u(x)$, BCs/ICs or parameters. However, the key difference stems in the fact that these other methods need to train on known solutions $s(x)$ for each of the (random-)sampled $u(x)$ functions, computed from traditional numerical methods. In our approach, we directly solve the DEs, so no previous knowledge from any solutions is needed. In this way, the method discussed in the paper is a different and complementary approach to DeepONets or FNOs.
@@
Stiff DEs, e.g. DEs where there is a large separation of scales between different features of the solution, can be challenging to solve [10, 11]. Stiffness can manifest in various forms, including eigenvalue-based stiffness (where the system’s Jacobian has widely varying eigenvalues), multi-scale stiffness (involving both fast and slow time scales such as the flame equation (strictly speaking, the flame equation exhibits multiple types of stiffness) ), damping-induced stiffness (common in physical systems with strong damping such as the Van der Pol oscillator), nonlinear stiffness, interval-dependent stiffness, and transient stiffness [12]. In the cases we present, traditional numerical methods are capable of solving the equations. However, in the context of solving DE systems with PINNs, we present differential geometry techniques that significantly improve, or even unlock, our ability to solve them [13]. Moreover, these techniques are particularly relevant when there is no established numerical setup to solve the problem. This is the case of inverse problems. PINNs are a well-suited tool for solving such problems [6].
There are previous works where concepts of differential geometry have been used in the context of NNs. In [14] the authors explore and characterize the behavior of NNs during training through the theory of Riemannian geometry and curvature - they relate the output of the NN with the input transformations through the curvature equations. An analysis and comparison between the convergence properties of the metric of NNs (used for classification) and the metric under the Ricci-DeTurck flow is done in [15]. In [16, 17], the authors remove edges of low importance (prune) the computational graph using the definition of Ricci curvature. More concretely, some recent works have proven that differential geometry techniques can increase the performance and accuracy of PINNs [18, 19, 20].
@@
Here, we present two efficient algorithms to use in PINNs that are aimed at significantly improving the solution of inverse problems for non-linear, coupled and multi-scale DEs. The first technique is called Multi-Head (MH) training, where we train the network to learn the general space of all solutions (or latent space) to different variations of the same DEs, from which we can perform transfer learning to different regimes [21, 22, 23, 24, 25, 26]. The second one is a technique that we call Unimodular regularization (UR) of the latent space of solutions. It is based on the computation of the metric tensor, a concept from differential geometry, induced on the latent-space-hypersurface. This second technique facilitates (or even unlocks) the transfer learning process to extreme regimes by ensuring a more controlled response of the latent space to changes in the DEs [27]. The purpose of this article is to introduce these techniques and show their efficiency with several examples. The algorithms that we present here are of general use in solving DEs and inverse problems. The article is structured as follows: in the “Methods” section we introduce the methodology used; we describe the NN structure in “Single head” and “Multi-head” subsections, the general training procedure in “Training and Transferring” subsection, and UR in last subsection. The application of these techniques to different ordinary differential equation (ODE) systems and results are presented in the “Results” section, showing consistent improvement in the obtained solutions. We solve the Flame Equation, the Van der Pol Oscillator, and Einstein Field Equations in 5-dimensional Anti de Sitter spacetime in its corresponding subsections respectively. The conclusions are summarized in the “Conclusions” section.
2 Methods
2.1 Context
@@
Physics-Informed Neural Networks (PINNs), first introduced in the works of Dissanayake and Phan-Thien [28] and Lagaris et al. [29], have emerged as a powerful framework for solving partial and ordinary differential equations (PDEs/ODEs). Raissi et al. [13] popularized this approach by demonstrating its effectiveness on a variety of challenging physical problems, while other researchers have made significant advancements in applying neural networks to PDEs and ODEs, including notable examples such as [30, 31, 32]. In this paradigm, one typically trains a distinct neural network for each unknown function in the governing equations. These neural networks are typically implemented as multilayer perceptrons (MLPs), a class of models composed of stacked fully connected layers interleaved with nonlinear activation functions. The training process involves minimizing a loss function that encodes the squared residuals of these differential equations, thereby embedding the physics directly into the learning objective. For different approaches to embed the physics using machine learning techniques we refer the reader to the following references [33, 34, 35].
@@
While traditional numerical methods often outperform PINNs in terms of computational efficiency and accuracy, PINNs offer unique advantages. One key benefit is their ability to learn the parametric dependence of solutions with respect to varying equation forms or initial/boundary conditions (IC/BC). In contrast, traditional methods typically require rerunning the solver for each set of different IC/BC. With PINNs, however, it is possible to train the network to simultaneously accommodate multiple conditions, enabling rapid evaluation of solutions for various IC/BC values or even different, parametrically related equations. By training on diverse initial and boundary condition samples, our approach can recover a free form function within the region covered by the observed data. We caution, however, that accuracy is generally limited to the span of the conditions seen during training and may not extend reliably to entirely unobserved regions of the input space. However, in the third example of this work (subsection “Einstein Field Equations” in the “Results” section) the available initial and boundary condition samples sufficiently cover the region of interest, so the functional recovery remains accurate under this assumption.
Another notable advantage of using PINNs is their ability to leverage transfer learning. Transfer learning allows the model to generalize beyond the range of its training data, enabling the prediction of solutions in unseen regions or conditions. This technique has proven crucial for obtaining solutions in the stiffer regimes of certain differential equations [26]. Throughout this paper, we will leverage this concept within the so-called MH framework.
@@
The latent space of solutions can then be “projected down” to a specific solution for a given element of the family of the DE system by non-linearly combining the $d$ components. This projection is performed by another NN, known as the head. We will describe the multi-head setup in detail in the following subsections. For pedagogical clarity, we will first discuss a single-head model to explain its implementation within the PINNs framework. Subsequent subsections will show how this approach can be generalized to multiple heads and how the latent space can be learned to extrapolate solutions to regions where the body has not been trained— a process known as transfer learning. Finally, we will introduce a technique called Unimodular Regularization, which improves the performance of the MH approach when transferring solutions to stiff regimes of differential equations.
@@ 2.2 Single-head
Figure 1: Scheme of the single head approach Neural Network. We can see how this simplified case is the standard case for a deep neural network with multiple hidden layers and one single output. The output solutions are calculated as $\psi^{\text{NN}}=\texttt{head}_{W}\left(H\right)$ with $\psi\in\mathbb{R}^{(m\times 1)}$, $H\in\mathbb{R}^{(m\times d)}$ and $W\in\mathbb{R}^{(d\times 1)}$, being $m$ the batch size (number of points sampled in the independent variable $x^{1}$ meshed with sampling of $x^{2,3,\dots}$), and $d$ the number of neurons in the last layer of the head. The loss $L_{\text{DE}}$ is of the form in Eq. (3), i.e. it contains the residuals of the DEs for which the NN is solving - this is the main characteristic of PINNs. The red arrow represents the updating of the NN-parameters $w_{j}$ (weighs and biases) based on minimization of the loss function through gradient descent (represented by $\nabla_{w_{j}}L_{\text{DE}}$).
@@
The sum over $\mu$ represents the total over the different elements of the DE system, including also all sampled points along the independent variable $x^{0}$. Finally, the sum over the equations is the total of the residuals for each equation in our system of DEs. Note that the distinction between summing and computing the mean is irrelevant, as it merely scales the loss by a factor inversely proportional to the number of sampled points.
@@ 2.3 Multi-head
The goal is to find a specific solution $\psi_{\alpha}^{\text{NN}}(x^{\mu})$ for each problem by projecting $H$ into a reduced space, non-linearly combining its components through a head with parameters $W_{\alpha}$, where the index $\alpha$ labels different variations of the DE family. Each head $W_{\alpha}$ correspond to a specific DE variation and produces a loss $L_{\text{DE},\alpha}$ of the form in Eq. 3. We then define the total loss at each iteration as follows:
$L_{\text{tot}}=\sum_{\alpha=1}^{\#\text{heads}}L_{\text{DE},\alpha}$ (4)
@@
Figure 2: Scheme of the multi-head approach Neural Network. The index $\alpha$ in the inputs $x^{\mu}_{\alpha}$ denotes all the variations within the DE family - all of them are inputs to the body, but each specific element of the family is later solved by its corresponding head, giving solutions $\psi_{\alpha}(x^{\mu}_{\alpha})$ as output ($\alpha=1,\dots,h$). The total loss $L_{\text{tot}}$ is as defined in (4). The backpropagation of the derivatives $\nabla_{w_{j}}L_{\text{tot}}$ indicated by the red arrow updates all the weights and biases $w_{j}$ (both of the body and all heads).
@@ 2.4 Training and transferring
Once the latent space $H$ stabilizes, we freeze the parameters of the body, so they no longer change, even as the model continues training. We denote the frozen body as $H^{f}$.
@@
The process of TL can be used in two different situations. The first one is to obtain solutions for new elements of the DE family given by choices of parameters/IC/BC/forcing functions whose values fall within the range of training of the body $H^{f}$. The second is to find a specific solution to a variation within the DE family, even if that variation lies outside the values used in training. This is often the case when transferring to stiffer regimes, since it will be easier to train the system in the non-stiff regime. Next, we present a technique used during training and show that it improves results during the TL phase.
Figure 3: Scheme of the transfer learning procedure from a pre-trained body through the Multi-Head approach. The blue, diagonal lines in the body $H^{f}$indicate that it is frozen, i.e. its weights and biases do not change. Only the weights $\widetilde{W}$ of the new head are learnt in order to find the solution to a new variation of the DE system, as indicated by the red arrow. As explained in the text, this method can be used both for getting solutions inside or outside the range in which the model has been trained on.
One might be concerned about the scalability of the multi-head approach to a large number of variations of the ODEs (or when generalizing to high-dimensional PDEs). A straightforward grid-based method needs $\mathcal{O}(N^{D})$ models for a $D$-dimensional space with $N$ points per dimension, making it impractical for high $D$. Conversely, our MH technique functions akin to a Monte Carlo sampler:
  1. 1.
    Linear growth in heads. We select $N$ sample points in parameter space (e.g. stiffness parameter) and train one head per point, so the total number of heads grows as $\mathcal{O}(N)$ rather than $\mathcal{O}(N^{D})$. Monte Carlo convergence theory ensures the approximation error decays as $N^{-1/2}$, independently of $D$.
  2. 2.
    Cost absorbed by the body. For high-dimensional PDEs or bundle solutions, the main computational effort is sampling the expanded input space—this impacts the shared “body” during both forward and backward passes, not the number of heads itself.
  3. 3.
    Comparison to VAEs. Although a VAE’s encoder and decoder complexity scales with the latent dimension, VAEs still require dense input or parameter sampling for training. Our MH approach trades latent-space complexity for Monte Carlo–style coverage.
In conclusion, MH requires a head for each sampled variation but mitigates the curse of dimensionality by utilizing Monte Carlo scaling ($\mathcal{O}(N)$ heads for $N$ samples), with exponential costs restricted to input sampling within the shared body, not through increased heads. Although we do not have a rigorous proof of Monte Carlo–style scaling, it is a well-justified assumption that has been observed empirically in many of our works.